# High-dimensional analysis of ridge regression for non-identically distributed data with a variance profile

### v3

@@ 1.3 Publicly available source code
https://github.com/Issoudab/RidgeRegressionVarianceProfile
to implement the experiments carried out in this paper.
@@ 2.1 High-dimensional linear regression from the random matrix perspective
In particular, for such data, the study of the minimum norm least-square estimator and the double descent behavior of the predictive risk has been considered in [HMRT22, Bac24, BHX20, RMR21]. The analysis of the predictive risk of ridge regression using iid data with a general covariance structure has been studied in [DW18, Bac24], while previous works on the statistical analysis of ridge regression from the RMT perspective include [EK18, hD16] and [CD11, TV04] for applications in wireless communication.
@@ 2.2 Linear regression for independent but non-identically distributed data
While the statistical analysis of linear regression for iid data with a general covariance structure is very well understood, the literature on the study of the linear model for non-identically distributed predictors appears to be scarcer. A first analysis of maximum likelihood estimation in standard models (including linear regression) for independent but non-identically distributed data dates back to [Ber82]. More recent works [BBK+19, KBB+20], on statistical inference in linear regression in the so-called model-free framework, allow to consider the setting on non-identically distributed predictors. The assumption of independent and identically distributed (iid) data often fails in real-world scenarios. This limitation notably appears in fields such as epidemiology [ETP+22], finance [Gro21], neuroscience [TvWM+13], and climatology [ZNM+21]. The recognition of this issue has motivated significant research into the analysis of non-iid data. For example, [ZMP22] highlights the non-iid nature of data in studies regarding the effectiveness of COVID-19 vaccines. They demonstrate that interactions between data points can introduce bias into estimation results. Beyond the risk of bias, the non-iid structure of data can also invalidate certain statistical methods. Generalized cross-validation, for instance, yields poor predictions in such contexts. In response to this challenge, [LLS24, AZVP24] have proposed robust alternatives specifically designed to handle dependencies between data points. The study of dependent data has also garnered growing attention in other domains. In control theory, researchers have made strides in addressing dependencies [NWB+20, TZMP23, ZTPM24]. Similarly, the field of signal denoising has seen advancements with methods that explicitly account for data correlations [SN23, KSS23]. These developments underscore the importance of adapting traditional methodologies to the challenges posed by non-iid data across a broad spectrum of disciplines. The study of correlated data is a growing topic in high-dimensional statistics [ZJVM24, ZM24]. However, to the best of our knowledge, the high-dimensional analysis of the linear model using non-identically distributed data has not considered so far. Hence we propose to tackle this framework by endowing variance profile to the data.
@@ 2.3 The use of variance profile in RMT
RMT allows to describe the asymptotic distribution of the eigenvalues of large matrices with random entries, see e.g. [BS10]. In particular, the well-known Marchenko-Pastur theorem characterizes the limiting spectral distribution of the covariance matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$ for a data matrix $X_{n}=Z_{n}\Sigma^{1/2}$ with iid rows in the asymptotic setting $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$.
However, in many applications, such as photon imaging [SHDW14], network traffic analysis [BMG13], ecology [AGB+15, AT15], neurosciences [ASS15, ARS15] or genomics for microbiome studies [CZL19], the amount of variance in the observed data matrix may significantly change from one sample to the other, that is between the rows of $X_{n}$. The literature on statistical inference from high-dimension matrices with heteroscedasticity has thus recently been growing [BM20, BDF17, LDS18, UHZB16, ZCW22, JFL22]. Modeling data as a random matrix $X_{n}=\Upsilon_{n}\circ Z_{n}$ with a variance profile to handle the setting of non-iid data has also found applications in the analysis of the performances of wireless digital communication channels [HLN07]. In the RMT literature, Hermitian random matrices with centered entries but non-equal distribution are referred to as generalized Wigner matrices for which many asymptotic properties are now well understood. For example, for Hermitian random matrices with a variance profile that is doubly stochastic (namely its rows and columns elements sump up to one), bulk universality at optimal spectral resolution for local spectral statistics have been established in [EYY12] and they are shown to converge to those of a standard Wigner matrix (that is with iid sub-diagonal entries). The case of a generalized Wigner matrix with a variance profile that is not necessarily doubly stochastic has been studied in [AEK17a], and non-hermitian random matrices with a variance profile have been considered in [CHNR18, HLN07, HLN06] using the notion of deterministic equivalent that consists in approximating the spectral distribution of a random matrix by a deterministic function. A recent work [BHX23] considers the analysis of approximate message passing for statistical estimation problems involving a data matrix with a variance profile, and an application to ridge regression to characterize the non-asymptotic distribution of the ridge estimator is proposed.
@@
[…]
Then, we build upon results from [HLN07] to construct deterministic equivalents (when $\lim_{n\to+\infty}p/n\to c>0$) of the Stieltjes transforms
$g_{\hat{\mu}_{n}}(z)=\frac{1}{p}\operatorname{Tr}[(\widehat{\Sigma}_{n}-zI_{p})^{-1}],\quad\mbox{for}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$
and
$g_{\tilde{\mu}_{n}}(z)=\frac{1}{n}\operatorname{Tr}[(\widetilde{\Sigma}_{n}-zI_{n})^{-1}],\quad\mbox{for}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$
of the empirical eigenvalue distribution $\hat{\mu}_{n}$ of $\widehat{\Sigma}_{n}$, and the empirical eigenvalue distribution $\tilde{\mu}_{n}$ of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{\top}$ respectively. Then, using [HLN07][Theorem 2.5], the following result holds.
@@
Under Assumptions 1.2-1.5, the following limit holds true almost surely
$\lim_{n\to\infty,\;p/n\to c}\left(g_{\hat{\mu}_{n}}(z)-\frac{1}{p}\operatorname{Tr}[T_{p}(z)]\right)=0,\quad\mbox{for all}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$
$\lim_{n\to\infty,\;p/n\to c}\left(g_{\tilde{\mu}_{n}}(z)-\frac{1}{n}\operatorname{Tr}[\widetilde{T}_{n}(z)]\right)=0,\quad\mbox{for all}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+},$
where
$T_{p}(z)=\mathop{\rm diag}\nolimits(T_{p}^{(1)}(z),\ldots,T_{p}^{(p)}(z))\quad\mbox{and}\quad\widetilde{T}_{n}(z)=\mathop{\rm diag}\nolimits(\widetilde{T}_{n}^{(1)}(z),\ldots,\widetilde{T}_{n}^{(n)}(z)),$
are diagonal matrices of size $p\times p$ and $n\times n$ respectively, whose diagonal elements are the unique solutions of the deterministic system of $p+n$ equations
$T_{p}^{(j)}(z)$ $=$ $\frac{-1}{z(1+(1/n)\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(z)])}\quad\mbox{for}\quad 1\leq j\leq p,$ (2.1)
$\widetilde{T}_{n}^{(i)}(z)$ $=$ $\frac{-1}{z(1+(1/n)\operatorname{Tr}[D_{i}T_{p}(z)])}\quad\mbox{for}\quad 1\leq i\leq n,$ (2.2)
where
$\widetilde{D}_{j}=\mathop{\rm diag}\nolimits(\gamma^{2}_{1j},\ldots,\gamma^{2}_{nj})\quad\mbox{and}\quad D_{i}=\mathop{\rm diag}\nolimits(\gamma^{2}_{i1},\ldots,\gamma^{2}_{ip}).$
Moreover, $\frac{1}{p}\operatorname{Tr}[T_{p}(z)]$ and $\frac{1}{n}\operatorname{Tr}[\widetilde{T}_{n}(z)]$ are the Stieltjes transforms of probability measures denoted as $\nu_{n}$ and $\tilde{\nu}_{n}$ respectively.
The measures $\nu_{n}$ and $\tilde{\nu}_{n}$, defined in Proposition 2.1, are called the deterministic equivalents of the empirical eigenvalue distributions $\hat{\mu}_{n}$ and $\tilde{\mu}_{n}$ respectively. By a slight abuse of notation, we may also denote by $T_{p}(z)$ the vector of size $p$ whose entries are the coefficients $T_{p}^{(j)}(z)$ for $1\leq j\leq p$ solutions of the fixed point equations (2.1). Currently, a classical method to numerically approximate the value of $T_{p}(z)$ is to solve the nonlinear system of deterministic equations (2.1) written in a vector form that is referred to as the Dyson equation in [AEK17a, AEK19, AEK17b, AEK18]. Indeed, as stated in [AEK17b][Theorem 2.1], the vector $T_{p}(z)$ is known to be the unique solution of the Dyson equation
@@
that corresponds to equations (2.1) and (2.2) written in a vector form, where $1/v$ has to be understood as taking the inverse of the elements of the vector $v$ entrywise. Similarly, the vector $\widetilde{T}_{n}(z)=(\widetilde{T}_{n}^{(i)}(z))_{1\leq i\leq n}$ satisfies the Dyson equation
$\frac{1}{\widetilde{T}_{n}(z)}=-zI_{n}+\frac{1}{n}\Gamma_{n}\frac{1}{1+\frac{1}{n}\Gamma_{n}^{\top}\widetilde{T}_{n}(z)}.$
In this paper, we sometimes consider, as an illustrative example, the specific class of quasi doubly stochastic variance profiles defined in Equation 1.6. The fixed point equation (2.3) typically does not have an explicit expression, and one has to rely on numerical methods to solve it as done in Section 4. Nevertheless, when the variance profile is quasi doubly stochastic, the solution of the Dyson equation is a vector $T_{p}(z)$ having constant entries equal to $m_{p}(z)\in\mathbb{C}$ (or equivalently $T_{p}(z)=m_{p}(z)I_{p}$ is a scalar matrix) that satisfies
@@
$\frac{1}{\tilde{m}_{n}(z)}=-z+\frac{p/n}{1+\tilde{m}_{n}(z)},\quad\mbox{for all}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+}.$
@@ 3 Main results
In this section, we derive deterministic equivalents for the DOF and the predictive risk of ridge regression. We also obtain a deterministic equivalent of the predictive risk of minimum norm least square estimation when the ridge regularization parameter tends to zero. We compare these results to those that are already known in the standard setting of iid data, and we highlight the emergence of the double descent phenomenon for non-iid data. Throughout this section, it is supposed that Assumptions 1.2-1.5 hold true.
@@ 3.1 Degrees of freedom
In this section, we prove that the quantity $df_{1}(\lambda)$ introduced in Section 1.1 is a deterministic equivalent of the DOF $\widehat{df}_{1}(\lambda)$
@@
Under Assumptions 1.2-1.3, for any $\lambda>0$, the following holds
@@
with $df_{1}(\lambda)=\operatorname{Tr}[\Sigma_{n}(\Sigma_{n}+\kappa(\lambda))^{-1}]$, and
@@
When the variance profile is quasi doubly stochastic, one has that $\operatorname{Tr}[\widetilde{D}_{j}]=n$. Moreover, as remarked in Section 2.3, the matrix $\widetilde{T}_{n}(-\lambda)=\tilde{m}_{n}(-\lambda)I_{n}$ is scalar, implying that for $1\leq j\leq p$, $\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]=n\tilde{m}_{n}(-\lambda)$. Moreover, in this setting, $\Sigma_{n}=I_{p}$. Consequently, if $\Gamma_{n}$ is quasi doubly stochastic, one has that $\kappa(\lambda)=\frac{1}{\tilde{m}_{n}(-\lambda)}I_{p},$ and the deterministic equivalent of the DOF is
$df_{1}(\lambda)=\operatorname{Tr}\left[\Sigma_{n}\left(\Sigma_{n}+\kappa(\lambda)\right)^{-1}\right]=p\left(1+\frac{1}{\tilde{m}_{n}(-\lambda)}\right)^{-1}.$
The above equality corresponds to the formula of a deterministic equivalent of the DOF derived in [Bac24] for iid data when the entries of the features matrix $X_{n}$ are made of iid centered random variables with variances equal to $1$.

### v4

@@ 1.3 Publicly available source code
https://github.com/Issoudab/RidgeRegressionVarianceProfile to implement the experiments carried out in this paper.
@@ 2.1 High-dimensional linear regression from the random matrix perspective
In particular, for such data, the study of the minimum norm least-square estimator and the double descent behavior of the predictive risk has been considered in [HMRT22, Bac24, BHX20, RMR21]. The analysis of the predictive risk of ridge regression using iid data with a general covariance structure has been studied in [DW18, Bac24]. More recently, the analysis of the minimum-norm least-squares estimator has enabled a substantial extension of asymptotic risk characterizations to various learning paradigms. In particular, [SBS24, PDT24] study the generalization properties of minimum-norm solutions in high-dimensional regimes, providing precise risk characterization in the settings of out-of-distribution prediction and transfer learning, building upon a rich line of earlier research that investigated ridge regression from a RMT perspective including [EK18, hD16, LS23]. In parallel, closely related RMT techniques were developed in the wireless communications literature, where ridge-type estimators naturally arise in signal detection and channel estimation problems; see, for example, [CD11, TV04].
@@ 2.2 Linear regression for independent but non-identically distributed data
While the statistical analysis of linear regression for iid data with a general covariance structure is very well understood, the literature on the study of the linear model for non-identically distributed predictors appears to be scarcer. A first analysis of maximum likelihood estimation in standard models (including linear regression) for independent but non-identically distributed data dates back to [Ber82]. More recent works [BBK+19, KBB+20], on statistical inference in linear regression in the so-called model-free framework, allow to consider the setting on non-identically distributed predictors. The assumption of independent and identically distributed (iid) data often fails in real-world scenarios. This limitation notably appears in fields such as epidemiology [ETP+22], finance [Gro21], neuroscience [TvWM+13], and climatology [ZNM+21]. The recognition of this issue has motivated significant research into the analysis of non-iid data. For example, [ZMP22] highlights the non-iid nature of data in studies regarding the effectiveness of COVID-19 vaccines. They demonstrate that interactions between data points can introduce bias into estimation results. Beyond the risk of bias, the non-iid structure of data can also invalidate certain statistical methods. In response to the limitations of the iid setting, [LLS24, AZVP24] have recently proposed robust cross-validation schemes specifically tailored to settings with dependent observations, explicitly accounting for serial correlation and structural instability. More broadly, the problem of cross-validation under dependence has attracted growing attention in the statistical and econometric literature. In particular, [RR22] provide a theoretical analysis of cross-validation procedures for dependent data, highlighting the failure modes of standard random-split approaches and establishing conditions under which modified validation schemes remain consistent. Similarly, [YR25] study cross-validation in contexts, proposing dependence-aware strategies and yield more reliable risk estimates. The study of dependent data has also garnered growing attention in other domains. In control theory, researchers have made strides in addressing dependencies [NWB+20, TZMP23, ZTPM24]. Similarly, the field of signal denoising has seen advancements with methods that explicitly account for data correlations [SN23, KSS23], and the study of correlated data is a growing topic in high-dimensional statistics [ZJVM24, ZM24].
@@ 2.3 The use of variance profile in RMT
RMT allows to describe the asymptotic distribution of the eigenvalues of large matrices with random entries, see e.g. [BS10]. In particular, the well-known Marchenko-Pastur theorem characterizes the limiting spectral distribution of the covariance matrix $\widehat{\Sigma}_{n}=\frac{1}{n}X_{n}^{\top}X_{n}$ for a data matrix $X_{n}=Z_{n}\Sigma^{1/2}$ with iid rows in the asymptotic setting $\lim\limits_{n\to+\infty}\frac{p}{n}=c>0$. A large class of random matrix models considered in RMT falls within the unitarily invariant framework [Mec14], in which the distribution of the random matrix is invariant under unitary (or orthogonal, in the real-valued case) conjugation. This structural property has played a central role in the modern theoretical analysis of high-dimensional statistical estimators. In particular, unitarily invariant random matrix models have been extensively studied in the context of linear ridge regression [LS23, DW18, Bac24]
However, in many applications (such as photon imaging [SHDW14], network traffic analysis [BMG13], ecology [AGB+15, AT15], neuroscience [ASS15, ARS15], or genomics and microbiome studies [CZL19]) one must go beyond this framework. Indeed, in these contexts, the amount of variance in the observed data matrix may vary substantially across samples, that is, across the rows of $X_{n}$. As a consequence, the assumption of homoscedasticity underlying unitarily invariant models is no longer appropriate. The literature on statistical inference for high-dimensional random matrices with heteroscedastic structures has therefore grown rapidly in recent years [BM20, BDF17, LDS18, UHZB16, ZCW22, JFL22]. This line of work motivates the study of non–unitarily invariant models, commonly referred to as variance-profiled matrices. In this setting, the dataset is modeled as a random matrix of the form $X_{n}=\Upsilon_{n}\circ Z_{n},$ with a variance profile to handle the setting of non-iid data has also found applications in the analysis of the performances of wireless digital communication channels [HLN07]. Such matrices provide a flexible modeling framework for capturing heterogeneity in the data, in particular in the presence of missing or sparse observations by setting selected entries of the variance profile matrix to zero.
In the RMT literature, Hermitian random matrices with centered entries but non-equal distribution are referred to as generalized Wigner matrices for which many asymptotic properties are now well understood. For example, for Hermitian random matrices with a variance profile that is doubly stochastic (namely its rows and columns elements sump up to one), bulk universality at optimal spectral resolution for local spectral statistics have been established in [EYY12] and they are shown to converge to those of a standard Wigner matrix (that is with iid sub-diagonal entries). The case of a generalized Wigner matrix with a variance profile that is not necessarily doubly stochastic has been studied in [AEK17a], and non-hermitian random matrices with a variance profile have been considered in [CHNR18, HLN07, HLN06] using the notion of deterministic equivalent that consists in approximating the spectral distribution of a random matrix by a deterministic function. A recent work [BHX23] considers the analysis of approximate message passing for statistical estimation problems involving a data matrix with a variance profile, and an application to ridge regression to characterize the non-asymptotic distribution of the ridge estimator is proposed.
@@
Let $\mu$ be a probability measure supported on $\mathbb{R}$}. Then, its Stieltjes transform is defined as $g_{\mu}(z)=\int_{{\mathbb{R}}}\frac{1}{t-z}d\mu(t),\quad\mbox{for}\quad{\color[rgb]{0,0,0}z}\in\mathbb{C}^{+},$ where $\mathbb{C}^{+}=\{z\in\mathbb{C},\Im(z)>0\}$ and $\Im(\cdot)$ denotes the imaginary part of a complex number.
Then, we build upon results from [HLN07] to construct deterministic equivalents of the Stieltjes transforms (when $\lim_{n\to+\infty}p/n\to c>0$) of the empirical eigenvalue distribution $\hat{\mu}_{n}$ of $\widehat{\Sigma}_{n}$, and the empirical eigenvalue distribution $\tilde{\mu}_{n}$ of $\widetilde{\Sigma}_{n}=\frac{1}{n}X_{n}X_{n}^{\top}$ respectively.
$g_{\hat{\mu}_{n}}(z)=\frac{1}{p}\operatorname{Tr}[(\widehat{\Sigma}_{n}-zI_{p})^{-1}],\quad\mbox{and}\quad g_{\tilde{\mu}_{n}}(z)=\frac{1}{n}\operatorname{Tr}[(\widetilde{\Sigma}_{n}-zI_{n})^{-1}]$
for $z\in\mathbb{C}\setminus{\mathbb{R}}^{+}$. Then, [HLN07][Theorem 2.5] states that the following result holds.
@@
Suppose that Assumption 1.2 to Assumption 3.2 hold true, then the following limit holds true almost surely
$\lim_{n\to\infty,\;p/n\to c}\left(g_{\hat{\mu}_{n}}(z)-\frac{1}{p}\operatorname{Tr}[T_{p}(z)]\right)=0,\quad\mbox{and}\quad\lim_{n\to\infty,\;p/n\to c}\left(g_{\tilde{\mu}_{n}}(z)-\frac{1}{n}\operatorname{Tr}[\widetilde{T}_{n}(z)]\right)=0,$
for all $z\in\mathbb{C}\setminus{\mathbb{R}}^{+}$ with $T_{p}(z)=\mathop{\rm diag}\nolimits(T_{p}^{(1)}(z),...,T_{p}^{(p)}(z))$ and $\widetilde{T}_{n}(z)=\mathop{\rm diag}\nolimits(\widetilde{T}_{n}^{(1)}(z),...,\widetilde{T}_{n}^{(n)}(z))$ diagonal matrices of size $p\times p$ and $n\times n$ respectively, whose diagonal elements are the unique solutions of the deterministic system of $p+n$ equations
$T_{p}^{(j)}(z)=\frac{-1}{z(1+(1/n)\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(z)])}\mbox{, }\widetilde{T}_{n}^{(i)}(z)=\frac{-1}{z(1+(1/n)\operatorname{Tr}[D_{i}T_{p}(z)])},$ (2.1)
with $1\leq j\leq p$, $1\leq i\leq n$, $\widetilde{D}_{j}=\mathop{\rm diag}\nolimits(\gamma^{2}_{1j},\ldots,\gamma^{2}_{nj})\mbox{ and }D_{i}=\mathop{\rm diag}\nolimits(\gamma^{2}_{i1},\ldots,\gamma^{2}_{ip}).$ Moreover, $\frac{1}{p}\operatorname{Tr}[T_{p}(z)]$ and $\frac{1}{n}\operatorname{Tr}[\widetilde{T}_{n}(z)]$ are the Stieltjes transforms of probability measures denoted as $\nu_{p}$ and $\tilde{\nu}_{n}$ respectively.
The measures $\nu_{n}$ and $\tilde{\nu}_{n}$, defined in Proposition 2.1, are called the deterministic equivalents of the empirical eigenvalue distributions $\hat{\mu}_{n}$ and $\tilde{\mu}_{n}$ respectively. By a slight abuse of notation, we may also denote by $T_{p}(z)$ the vector of size $p$ whose entries are the coefficients $T_{p}^{(j)}(z)$ for $1\leq j\leq p$ solutions of the fixed point Equations Equation (2.1). Currently, a classical method to numerically approximate the value of $T_{p}(z)$ is to solve the nonlinear system of deterministic Equations Equation (2.1) written in a vector form that is referred to as the Dyson equation in [AEK17a, AEK19, AEK17b, AEK18]. Indeed, as stated in [AEK17b][Theorem 2.1], the vector $T_{p}(z)$ is known to be the unique solution of the Dyson equation
@@
that corresponds to Equation Equation (2.1) written in a vector form, where $1/v$ has to be understood as taking the inverse of the elements of the vector $v$ entrywise. Similarly, the vector $\widetilde{T}_{n}(z)=(\widetilde{T}_{n}^{(i)}(z))_{1\leq i\leq n}$ satisfies the Dyson equation $\frac{1}{\widetilde{T}_{n}(z)}=-zI_{n}+\frac{1}{n}\Gamma_{n}\frac{1}{1+\frac{1}{n}\Gamma_{n}^{\top}\widetilde{T}_{n}(z)}.$ In this paper, we sometimes consider, as an illustrative example, the specific class of quasi doubly stochastic variance profiles defined in Equation (1.7). The fixed point equation Equation (2.2) typically does not have an explicit expression, and one has to rely on numerical methods to solve it as done in Section 4. Nevertheless, when the variance profile is quasi doubly stochastic, the solution of the Dyson equation is a vector $T_{p}(z)$ having constant entries equal to $m_{p}(z)\in\mathbb{C}$ (or equivalently $T_{p}(z)=m_{p}(z)I_{p}$ is a scalar matrix) that satisfies
@@
$\frac{1}{\tilde{m}_{n}(z)}=-z+\frac{p/n}{1+\tilde{m}_{n}(z)},\quad\mbox{for all}\quad z\in\mathbb{C}\setminus{\mathbb{R}}^{+}.$ (2.4)
@@ 3 Main results
In this section, we derive deterministic equivalents for the DOF, the GCV criterion and the predictive risk of ridge regression. We also obtain a deterministic equivalent of the predictive risk of minimum norm least square estimation when the ridge regularization parameter tends to zero. We compare these results to those that are already known in the standard setting of iid data, and we highlight the emergence of the double descent phenomenon for non-iid data. In the literature dealing with iid data, these equivalents are typically used to identify regimes of good and bad prediction, and to design tuning rules for $\lambda$. In the variance-profile setting, a similar analysis can be carried out but the conclusions naturally split into:
  • (a)
    Universal consequences that hold across all variance profiles, such as the optimal choice of $\lambda_{\ast}$ in Corollary 3.1;
  • (b)
    Profile-dependent consequences, such as the shape of the ridgeless risk as a function of $c=p/n$, which can significantly deviate from classical double descent.
The remainder of Section 3 establishes the deterministic equivalents needed to reach such conclusions.
@@ 3.1 Degrees of freedom
In this section, we prove that the quantity $df_{1}(\lambda)$ introduced in Section 1.1 is a deterministic equivalent of the DOF $\widehat{df}_{1}(\lambda)$ .
@@
Suppose that Assumptions 1.2 and 1.3 hold true, then for any $\lambda>0$, one has that :
@@
with $df_{1}(\lambda)=\frac{1}{p}\operatorname{Tr}[\Sigma_{n}(\Sigma_{n}+\kappa(\lambda))^{-1}]$, and
@@
When the variance profile is quasi doubly stochastic, one has that $\operatorname{Tr}[\widetilde{D}_{j}]=n$. Moreover, as remarked in Section 2.3, the matrix $\widetilde{T}_{n}(-\lambda)=\tilde{m}_{n}(-\lambda)I_{n}$ is scalar, implying that for $1\leq j\leq p$, $\operatorname{Tr}[\widetilde{D}_{j}\widetilde{T}_{n}(-\lambda)]=n\tilde{m}_{n}(-\lambda)$. Moreover, in this setting, $\Sigma_{n}=I_{p}$. Consequently, if $\Gamma_{n}$ is quasi doubly stochastic, one has that $\kappa(\lambda)=\frac{1}{\tilde{m}_{n}(-\lambda)}I_{p},$ and the deterministic equivalent of the DOF is $df_{1}(\lambda)=\frac{1}{p}\operatorname{Tr}\left[\Sigma_{n}\left(\Sigma_{n}+\kappa(\lambda)\right)^{-1}\right]=\left(1+\frac{1}{\tilde{m}_{n}(-\lambda)}\right)^{-1}.$ The above equality corresponds to the formula of a deterministic equivalent of the DOF derived in [Bac24] for iid data when the entries of the features matrix $X_{n}$ are made of iid centered random variables with variances equal to $1$.